Solve

3. yy = 2y2 In(x) + 2xy2 – In(x) –
Answer #1

Solution: solve the differential equation yy = 242 loca) + guy? - b(A) - x - (lon ta) - yyl= gya (lola) + x + x) - gyl = (lo=> 11 Seondu [xdx 14u --> du (somes for -- d (en) du fordu | 4 + Inful = (Anne) (sc) xx I de + С + I lolul = x.lo (n) =) fals